# Blow-up for a Nonlocal Diffusion Equation with Time Regularly Varying Nonlinearity and Forcing

### v2

@@ 1. Introduction and main results
The structure of the article is as follows. In Section 2, we introduce the notation used throughout the paper and present several auxiliary results and estimates. Section 3 is devoted to the study of the unforced problem (1.3), where we provide the proof of Theorem 1.1. In Section 4, we address the main problem (1.1) and establish Theorems 1.2 and 1.3. Concluding remarks and directions for future research are given in Section 5. Finally, Appendix A contains a brief overview of regularly varying functions together with several useful estimates employed in our analysis.
@@ 3. The unforced problem
In this section, we provide the proof of Theorem 1.1, which addresses the unforced problem (1.3).
@@
    Suppose that $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}.$ According to [15, Theorem 6], it suffices to show that (1.5) holds for some $t_{0}>0$. By Lemma 2.3, we have
@@
    Moreover, by (A.14), there exists a slowly varying function $\ell$ such that, for all sufficiently large $t>0$,
@@
    Consequently, for $t>0$ large enough,
    $\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\geq\;\int\limits_{t/2}^{t}\mathsf{h}(\tau)\,d\tau\;\gtrsim\;t^{\gamma+1}.$ (3.3)
    Inserting (3.3) into (3.1) yields
    $(p-1)\,\|e^{t\mathscr{L}}u_{0}\|_{\infty}^{p-1}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\gtrsim\;t^{\gamma+1-\frac{N(p-1)}{2\mathsf{s}}}.$ (3.4)
    Since $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}$, we have $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}>0.$ Therefore, one can choose $t_{0}>0$ sufficiently large so that (1.5) holds. This completes the proof of the first part of Theorem 1.1.
@@
    Assume now that $p>1+\frac{2\mathsf{s}(\gamma+1)}{N}.$ We will prove that condition (1.4) holds, which guarantees the global existence of solutions. For $t_{0}>0$ (to be chosen sufficiently large later), applying Lemma 2.3 together with (A.14), we obtain
@@
    where $\ell_{1}$ is a slowly varying function provided by Theorem A.4. Since $\gamma-\frac{N(p-1)}{2\mathsf{s}}<-1,$ Theorem A.3 yields
@@
    Finally, since $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}<0$, the desired result follows from [15, Theorem 6], by first choosing $t_{0}>0$ sufficiently large and then taking $\|v_{0}\|_{\infty}$ small enough.

### v3

# Fujita Phenomenon for Mixed Local–Nonlocal Diffusion Equations

@@ 1. Introduction and main results
    Although the class $\mathcal{M}(\gamma)$ allows for oscillatory and non-regular behavior, the blow-up versus global existence threshold is unchanged. This highlights the robustness of the Fujita phenomenon with respect to temporal irregularities in the coefficient $h$, provided its logarithmic growth rate is controlled.
  3. (iii)
@@
The structure of the article is as follows. In Section 2, we introduce the notation used throughout the paper and present several auxiliary results and estimates. Section 3 is devoted to the study of the unforced problem (1.3), where we provide the proof of Theorem 1.2. In Section 4, we address the main problem (1.1) and establish Theorems 1.4 and 1.6. Concluding remarks and directions for future research are given in Section 5. Finally, Appendix A contains a brief overview of regularly varying functions together with several useful estimates employed in our analysis.
@@ 3. The unforced problem
In this section, we present the proof of Theorem 1.2, which concerns the unforced problem (1.3). Our approach follows the general strategy developed in [15, Theorem 6]. The proof consists in verifying the sufficient criteria for blow-up and global existence stated therein. Special attention is required because the coefficient $h$ belongs to the class $\mathcal{M}(\gamma)$, which is strictly larger than the class of regularly varying functions.
@@
    Assume that $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}$. According to [15, Theorem 6], it is enough to show that condition (1.5) holds for some $t_{0}>0$. By Lemma 2.3, we obtain
@@
    Furthermore, by (A.14), there exists a slowly varying function $\ell$ such that, for all sufficiently large $t>0$,
@@
    Hence, for $t$ large enough,
    $\begin{split}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;&\geq\;\int\limits_{t/2}^{t}\mathsf{h}(\tau)\,d\tau\\ \;&\gtrsim\;\int\limits_{t/2}^{t}\tau^{\gamma}\,\ell(\tau)\,d\tau\\ \;&\gtrsim\;t^{\gamma+1}\int\limits_{1/2}^{1}r^{\gamma}\,\ell(tr)\,dr.\end{split}$ (3.3)
    By Theorem A.3, as $t\to\infty$ we have
    $\int\limits_{1/2}^{1}r^{\gamma}\,\ell(tr)\,dr\;\sim\;\ell(t).$ (3.4)
    Combining (3.4) and (3.3) with (3.1), we deduce
    $(p-1)\,\|e^{t\mathscr{L}}u_{0}\|_{\infty}^{p-1}\int\limits_{0}^{t}\mathsf{h}(\tau)\,d\tau\;\gtrsim\;t^{\gamma+1-\frac{N(p-1)}{2\mathsf{s}}}\,\ell(t).$ (3.5)
    Since $p<1+\frac{2\mathsf{s}(\gamma+1)}{N}$, it follows that
    $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}>0.$
    Therefore, in view of (A.6), one can choose $t_{0}>0$ sufficiently large so that (1.5) is satisfied. This concludes the proof of the first part of Theorem 1.2.
@@
    We now assume that $p>1+\frac{2\mathsf{s}(\gamma+1)}{N}$. We will show that condition (1.4) holds, which ensures the global existence of solutions. Let $t_{0}>0$ (to be fixed sufficiently large later). By Lemma 2.3 and (A.14), we obtain
@@
    where $\ell_{1}$ is a slowly varying function given by Theorem A.11. Since $\gamma-\frac{N(p-1)}{2\mathsf{s}}<-1$ and the function
    $\tau\longmapsto\ell_{1}(\tau)\,\tau^{\gamma-\frac{N(p-1)}{2\mathsf{s}}}$
    belongs to the class $\mathcal{RV}_{\gamma-\frac{N(p-1)}{2\mathsf{s}}}$, Theorem A.7 yields
@@
    Finally, since $\gamma+1-\frac{N(p-1)}{2\mathsf{s}}<0$, the conclusion follows from [15, Theorem 6] by first choosing $t_{0}>0$ sufficiently large and then taking $\|v_{0}\|_{\infty}$ small enough.